# The Energy Cost of Artificial Intelligence Lifecycle in Communication Networks

### v3

@@ IV Energy cost of Data Collection
In this section, we derive the energy cost of data collection in an AI-powered wireless system. To calculate this cost, we first determine the number of bits that need to be transmitted from an device to the server via wireless technology (uplink). Then, we calculate the energy consumption of both transmitting and receiving the data, based on the given transmission and reception power and the corresponding rate. Finally, we derive the total energy consumption for collecting data from the device. It is important to note that the energy consumption of wired communication is considered negligible in comparison to the wireless part [31]. Therefore, this work focuses solely on the communication from the wireless device to the access point (AP) via wireless technologies. The assumption is that all data is collected via a single link; however, a summation can be added for multiple links. The modular and extensible open source calculator can be extended to include specific communication segments, topologies and technologies employing various resource allocation and optimization techniques.
IV-A OSI Overhead
@@
$B_{\mathrm{T}}$ $\leavevmode\nobreak\ [b]=\lceil B_{\mathrm{T,L_{\mathrm{OSI}}}}\prod_{l=1}^{L_{\mathrm{OSI}}}\underbrace{\left(RR_{\mathrm{DP},l}(1+OH_{\mathrm{DP},l})\right.}_{\text{contribution from DP}}$
@@
$+\underbrace{\left.RR_{\mathrm{CP},l}\gamma_{l}OH_{\mathrm{CP},l}\right)}_{\text{contribution from CP}}\rceil,\leavevmode\nobreak\ \mathrm{where}\leavevmode\nobreak\ L_{\mathrm{OSI}}=7,$ (2)
@@
where $B_{\mathrm{T,L_{OSI}}}$ represents the application layer data (i.e. telemetry) sent from the device, measured in bits. It is calculated as the product of the bit precision ($f$), the number of samples ($N_{\mathrm{S}}$) and the sample size ($I_{\mathrm{S}}$) in floating-point format, i.e., $B_{\mathrm{T,L_{\mathrm{OSI}}}}=fN_{\mathrm{S}}I_{\mathrm{S}}$. Typically, $f$ equals $32$ bits for single precision (float) or 64 bits for double precision format. To be able to send/receive data, in each of the seven OSI layers, represented by the product in Eq. 2, a number of data and control plane operations take place as expressed in the respective terms. The application data, $B_{\mathrm{T,L_{OSI}}}$, is processed and is added some overhead $OH_{\mathrm{DP},l}$, in some layers it is encrypted (i.e. application, presentation and/or transport layers) and in some layers it may also require retransmissions $RR_{\mathrm{DP},l}$. Moreover, some layers, such as the network layer, incur additional overhead required from signaling to maintain routing tables. This is reflected in the second term of Eq. 2 through the respective CP overheads $OH_{\mathrm{CP},l}$, retransmissions $RR_{\mathrm{CP},l}$ and the $\gamma_{l}$ scaling factor that adjusts for the relative contribution of the CP overhead compared to the DP counterpart at the $l$-th layer. The scaling factor can be also used to compensate for other specific overheads unforeseen by the considered configuration in this paper.
@@
The eCAL calculator implements the computation from Eq. 2 and enables the configuration of the application data size, retransmission rates, overheads, and scaling factor. Furthermore, it enables plugging custom protocol implementations for each of the layers, based also on the existing literature [32, 33, 34], so that the community can also instantiate eCAL for a specific well-defined set of protocols in addition to the approach adopted in this paper.
@@
To illustrate the behavior of Eq. 2, we first consider a scenario where a single device that transmits $N_{\mathrm{S}}$ of double-precision samples to the application accompanied by DP and CP overheads while altering the retransmission rate. In particular, we consider that each layer in the OSI model introduces a fixed overhead percentage for the DP and CP, with $10$% and $5$%, respectively. Furthermore, the relative contribution of CP overhead is scaled by a factor of $0.5$, i.e., $\gamma_{l}=0.5$. The retransmission rates for both DP and CP are assumed to be equal and vary from $1$ to $2$ in increments of $0.04$. This is shown in Fig. 2. The results show the retransmission rate has a significant impact on total transmitting bits. For example, when $N_{\mathrm{S}}=256$ samples, there is a growth of $2.3$, $17$, and $128$ times with retransmission rates equal to $1.125$, $1.5$, and $2$, respectively, compared to scenarios with no retransmission.
@@
IV-B Energy Estimation
@@
To estimate the energy consumption of the data collection component ($E_{\mathrm{DC}}$), which is the sum of the energy of transmitting ($E_{\mathrm{T}}$) and receiving the data ($E_{\mathrm{R}}$) introduced in the physical layer (PHY), as well as the energy consumption of computation in the DP and CP across the various layers of OSI ($E_{\mathrm{C}}$). We propose the following analytical expression:
$E_{\mathrm{DC}}$ $=\underbrace{\frac{P_{\mathrm{T}}[W]}{R_{\mathrm{T}}[b/s]}B_{\mathrm{T}}[b]}_{\text{PHY layer transmitting energy ($E_{\mathrm{T}}$)}}+\underbrace{\frac{P_{\mathrm{R}}[W]}{R_{\mathrm{R}}[b/s]}B_{\mathrm{T}}[b]}_{\text{PHY layer receiving energy ($E_{\mathrm{R}}$)}}$ (3)
$+\sum^{L_{\mathrm{OSI}}}_{l=2}(\underbrace{B_{\mathrm{T}_{l}}\leavevmode\nobreak\ [b]N_{\mathrm{dev,cycle},l}\leavevmode\nobreak\ [c/b]P_{\mathrm{dev,cycle}}\leavevmode\nobreak\ [J/c]}_{\text{from APP to MAC layer computational energy at the device}}$
$+\underbrace{B_{\mathrm{T}_{l}}\leavevmode\nobreak\ [b]N_{\mathrm{gateway,cycle},l}\leavevmode\nobreak\ [c/b]P_{\mathrm{gateway,cycle}}\leavevmode\nobreak\ [J/c]}_{\text{from MAC to APP layer computational energy at the gateway}})$
$=E_{\mathrm{T}}+E_{\mathrm{R}}+E_{\mathrm{C}},$
@@
where $P_{\mathrm{T}}$ and $R_{\mathrm{T}}$ represent the transmitting power and the data rate, respectively, while $P_{\mathrm{R}}$ and $R_{\mathrm{R}}$ are the receiving power and data rate. Furthermore, $B_{\mathrm{T},l}$ denotes the total number of the bits that need to be processed at the $l$-th layer in the OSI model. Additionally, we define the computational parameters at the device as $N_{\mathrm{dev,cycle},l}$ representing the number of processing cycles per bit at the $l$-th layer, and $P_{\mathrm{dev,cycle}}$, denoting the power consumption per cycle. Similarly, at the gateway, these parameters are defined as $N_{\mathrm{gateway,cycle},l}$ and $P_{\mathrm{gateway,cycle}}$.
@@
Figure 3: Energy consumption of data collection ($E_{\mathrm{DC}}$) (log scale) with its components and its components, including transmission energy ($E_{\mathrm{T}}$), receiving energy ($E_{\mathrm{R}}$), and computational energy ($E_{\mathrm{C}}$) with $N_{\mathrm{S}}=256$, $RR_{\mathrm{DP},l}=RR_{\mathrm{CP},l}=1$, $OH_{\mathrm{DP},l}=10\%$, $OH_{\mathrm{CP},l}=5\%$, and $\gamma_{l}=0.5$.
@@
Computing exact values for $E_{\mathrm{DC}}$ for a broad range of $B_{\mathrm{T}_{l}}$, $N_{\mathrm{\_,cycle},l}$, and $P_{\mathrm{\_,cycle}}$ is challenging due to the overall complexity of computing systems. For instance, $N_{\mathrm{\_,cycle,}l}$ varies across microprocessors when processing the same amount of data, and it does not have a linear dependency with data size [35]. Therefore, the values for $B_{\mathrm{T}_{l}}$ and $N_{\mathrm{\_,cycle}}$ cannot be chosen arbitrarily. Next, $P_{\mathrm{\_,cycle}}$ also depends on the microprocessor. However, in Fig. 3, we illustrate the energy consumption of $E_{\mathrm{DC}}$ with a selected set of values to provide an intuition on the contribution from each term. More specifically, we analyze a scenario where $N_{\mathrm{S}}=256$, and we consider two sets of parameters for computing transmitting and receiving energy, i.e. $E_{\mathrm{T}}$ and $E_{\mathrm{R}}$, alongside two different settings for calculating the computational energy consumption ($E_{\mathrm{C}}$). The results show that while transmission energy ($E_{\mathrm{T}}$) remains the dominant contributor in most cases, the computational energy consumption ($E_{\mathrm{C}}$) is not negligible and, in some configurations, constitutes a significant portion of the total energy consumption of data collection. This is particularly the case when the hardware is less optimized for the tasks [35], leading to higher cycles per bit ($N_{\mathrm{\_,cycle},l}$) and power per cycle ($P_{\mathrm{\_,cycle}}$). For example, when $N_{\mathrm{\_,cycle}}=100$ and $P_{\mathrm{\_,cycle}}=2\times 10^{-10}$, $E_{\mathrm{C}}$ is $7.55\times 10^{-3}\leavevmode\nobreak\ [J]$, which represents approximately $10\%$ and $95\%$ of the total energy cost in two different PHY configurations. On the other hand, when $N_{\mathrm{\_,cycle},l}=50$ and $P_{\mathrm{\_,cycle}}=10^{-10}$, $E_{\mathrm{C}}$ is $6.61\times 10^{-3}\leavevmode\nobreak\ [J]$, contributing to approximately $5\%$ and $90\%$ of the total energy consumption in data collection component.
@@ V Energy Cost of Data Preprocessing
At the server side, data is first stored on a hard drive, subsequently preprocessed, and finally used for the ML model development or inference. This section focuses on preprocessing, while inference is addressed in Section VII. Storage is generally energy-efficient, with relatively negligible energy costs. However, it can become significant when handling large volumes of incoming data or when utilizing virtualized hosted storage with encryption in an edge-cloud environment [36]. In the following case study, we consider the dataset with $N_{\mathrm{S}}$ samples collected via the wireless network, as depicted in Fig. 1 and discussed in Section IV. We also consider data preprocessing consisting of data cleaning, and data transformation steps, while leaving other more complex approaches, such as feature engineering, for further extensions of the study and simulator.
In the data cleaning step, the task of preprocessing is to identify and remove invalid samples from a list or array, which incurs zero floating-point operations (FLOPs [^3] FLOPs are often use to enable theoretical estimation of computational complexity. Alternative approaches would involve measurements through interfaces such as PAPI or tools such as Scaphandre when measuring directly energy., as no floating-point additions or multiplications are performed, only comparisons and memory operations. During the data transformation step, preprocessing involves determining the complexity of data transformation ($M_{\mathrm{DS}}$). We provide three examples, i.e., min-max scaling, normalization, and Gramian Angular Difference Field (GADF) process which represent typical approaches for data transformation.
@@
Standard score, also known as z-score normalization, represents a more computationally intensive preprocessing technique compared to min-max scaling. The process involves three distinct computational steps. First, calculating the population mean requires $N_{\mathrm{S}}I_{\mathrm{S}}$ FLOPs, as we sum all values and divide by the total number of samples. Second, computing the population standard deviation demands $13N_{\mathrm{S}}I_{\mathrm{S}}+$ 1 FLOPs, which encompasses calculating squared differences from the mean, their sum, and the final square root operation. Finally, the standardization transformation itself necessitates $2N_{\mathrm{S}}I_{\mathrm{S}}$ FLOPs, as each data point must be centered by subtracting the mean and scaled by dividing by the standard deviation. The total computational cost for normalization, expressed in FLOPs ($M^{\mathrm{norm}}_{\mathrm{DS}}$), is therefore:
@@
To showcase a more computationally expensive preprocessing operation, we implemented a calculator to compute the energy cost of the GADF transformation [37]. This transformation converts time series data into an image representation that captures temporal correlations between each pair of values in the time series. The GADF computation consists of several sequential steps, each contributing to the overall computational complexity and, consequently, the energy consumption. The first step requires scaling the time series data using min-max scaling, which ensures values fall within the [-1, 1] interval. Following the scaling, the data undergoes conversion to polar coordinates, where the radius is calculated using the time stamp values, and the angular coordinates are computed through an inverse cosine operation. The final step involves constructing the GADF matrix through pairwise angle differences, resulting in an $I_{\mathrm{S}}\times I_{\mathrm{S}}$ matrix where $I_{\mathrm{S}}$ represents the length of the input time series.
The computational complexity of GADF can be broken down into its constituent operations. The initial min-max scaling requires $12N_{\mathrm{S}}I_{\mathrm{S}}+$ 1 FLOPs as previously discussed, however in these case we have to multiply the number of samples with the length of the input time series ($I_{\mathrm{S}}$) as we are dealing with a series of values in a single sample increasing the FLOPs. The polar coordinate transformation and the construction of the GADF matrix itself, requires $(5I_{\mathrm{S}}+I_{\mathrm{S}}^{2})N_{\mathrm{S}}$ FLOPs . Therefore, the total number of FLOPs ($M^{\mathrm{GADF}}_{\mathrm{DS}}$) for the GADF transformation can be expressed as:
@@
This quadratic computational complexity makes GADF significantly more energy-intensive compared to simpler preprocessing operations like standardization or min-max scaling, particularly for longer time series.
As a result, the total number of FLOPs for preprocessing can be expressed as:
@@
With the computing complexity in FLOPs formalized, we can obtain total energy consumption ($E_{\mathrm{pre}}$) of the preprocessing ($E_{\mathrm{pre}}$) by utilizing the power consumption of the processing unit ($P_{\mathrm{pre}}$) and the executing time ($T_{\mathrm{pre}}$). Thus, it is expressed as follows:
@@
[…]

### v4

@@ IV Energy cost of Data Collection
In this section, we derive the energy cost of data collection in an AI-powered wireless system. To calculate this cost, we first determine the number of bits that need to be transmitted from a device to the server via wireless technology (uplink). Then, we calculate the energy consumption of transmitting and receiving the data, based on the given transmission and reception power and the corresponding data rate. Finally, we derive the total energy consumption for collecting data from a device. The assumption is that all data is collected via a single link; however, a summation can be added to generalize data collection to multiple links with different parameters. In this regard, the modular and extensible open source calculator can be extended to include specific communication segments, topologies, and technologies employing various resource allocation and optimization techniques.
IV-A OSI Overhead Estimation Method in eCAL
@@
$B_{\mathrm{T}}$ $\penalty 10000\ [b]=\lceil B_{\mathrm{T,L_{\mathrm{OSI}}}}\prod_{l=1}^{L_{\mathrm{OSI}}}\underbrace{\left(RR_{\mathrm{DP},l}(1+OH_{\mathrm{DP},l})\right.}_{\text{contribution from DP}}$
@@
$+\underbrace{\left.RR_{\mathrm{CP},l}\gamma_{l}OH_{\mathrm{CP},l}\right)}_{\text{contribution from CP}}\rceil,\penalty 10000\ \mathrm{where}\penalty 10000\ L_{\mathrm{OSI}}=7,$ (2)
@@
where $B_{\mathrm{T,L_{OSI}}}$ represents the application layer data (e.g. telemetry) sent from the device, measured in bits. It is calculated as the product of the bit precision ($f$), the number of samples ($N_{\mathrm{S}}$) and the sample size ($I_{\mathrm{S}}$) in floating-point format, i.e., $B_{\mathrm{T,L_{\mathrm{OSI}}}}=fN_{\mathrm{S}}I_{\mathrm{S}}$. Typically, $f$ equals $32$ bits for single precision (float) or 64 bits for double precision format. To be able to send/receive data, in each of the seven OSI layers, represented by the product in Eq. (2), a number of data and control plane operations take place as expressed in the respective terms. The application data, $B_{\mathrm{T,L_{OSI}}}$, is processed and is added some overhead $OH_{\mathrm{DP},l}$, in some layers it is encrypted (i.e. application, presentation and/or transport layers) and in some layers it may also require retransmissions $RR_{\mathrm{DP},l}$. Moreover, some layers, such as the network layer, incur additional overhead required from signaling to maintain routing tables. This is reflected in the second term of Eq. (2) through the respective CP overheads $OH_{\mathrm{CP},l}$, retransmissions $RR_{\mathrm{CP},l}$ and the $\gamma_{l}$ scaling factor that adjusts for the relative contribution of the CP overhead compared to the DP counterpart at the $l$-th layer. The scaling factor can also be used to compensate for other specific overheads unforeseen by the considered configuration in this paper, such as the fiber-based wired communication backhaul.
Numerical Validation
@@
The eCAL calculator implements the computation from Eq. (2) and enables the configuration of the application data size, retransmission rates, overheads, and scaling factor. Furthermore, it enables plugging custom protocol implementations for each of the layers, based also on the existing literature [30, 31, 32], so that the community can also instantiate eCAL for a specific well-defined set of protocols in addition to the approach adopted in this paper.
@@
To illustrate the behavior of Eq. (2), we first consider a scenario where a single device transmits $N_{\mathrm{S}}$ of double-precision samples to the application accompanied by DP and CP overheads while altering the retransmission rate. In particular, we consider that each layer in the OSI model introduces a fixed overhead percentage for the DP and CP, with 1010% and 55%, respectively. Furthermore, the relative contribution of CP overhead is scaled by a factor of $0.5$, i.e., $\gamma_{l}=0.5$. The retransmission rates for both DP and CP are assumed to be equal and vary from $1$ to $2$ in increments of $0.04$. This is shown in Fig. 2. The results show the retransmission rate has a significant impact on total transmitting bits. For example, when $N_{\mathrm{S}}=256$ samples, there is a growth of $2.3$, $17$, and $128$ times with retransmission rates equal to $1.125$, $1.5$, and $2$, respectively, compared to scenarios with no retransmission.
@@
IV-B Application Layer Centered Energy Estimation for the Data Collection Component of eCAL
@@
To estimate the energy consumption of the data collection component ($E_{\mathrm{DC}}$), which is the sum of the energy of transmitting ($E_{\mathrm{T}}$) and receiving the data ($E_{\mathrm{R}}$) introduced in the physical layer (PHY), as well as the energy consumption of computation in the DP and CP across the various layers of OSI ($E_{\mathrm{C}}$). We consider a network comprising $N_{\mathrm{L}}$ independent and direct links between nodes, which covers all the network topologies, including multi-access and multi-hop links. We propose the following analytical expression for the $p$-th links, which includes:
$E_{\mathrm{DC},p}$ $=$ (3)
[…]
@@
where $P_{\mathrm{T},p}$ and $R_{\mathrm{T},p}$ represent the transmitting power and the data rate, respectively, while $P_{\mathrm{R},p}$ and $R_{\mathrm{R},p}$ are the receiving power and data rate. Furthermore, $B_{\mathrm{T},l}$ denotes the total number of bits that need to be processed at the $l$-th layer in the OSI model. Additionally, we define the computational parameters at the device as $N_{\mathrm{dev,cycle},l,p}$ representing the number of processing cycles per bit at the $l$-th layer, and $P_{\mathrm{dev,cycle}}$, denoting the power consumption per cycle. Similarly, at the gateway, these parameters are defined as $N_{\mathrm{gateway,cycle},l}$ and $P_{\mathrm{gateway,cycle}}$. As a result, we can calculate the total energy cost of the data collection as the sum of all $N_{\mathrm{L}}$ links, which is expressed as:
$E_{\mathrm{DC}}=\sum^{N_{\mathrm{L}}}_{p=1}E_{\mathrm{DC},p}.$ (4)
@@
Figure 3: Energy consumption of data collection ($E_{\mathrm{DC}}$) (log scale) with its components and configurations, including transmission energy ($E_{\mathrm{T}}$), receiving energy ($E_{\mathrm{R}}$), and computational energy ($E_{\mathrm{C}}$) with $N_{\mathrm{S}}=256$, $RR_{\mathrm{DP},l}=RR_{\mathrm{CP},l}=1$, $OH_{\mathrm{DP},l}=10\%$, $OH_{\mathrm{CP},l}=5\%$, and $\gamma_{l}=0.5$.
Numerical Validation
@@
Computing exact values for $E_{\mathrm{DC}}$ for a broad range of $B_{\mathrm{T}_{l}}$, $N_{\mathrm{\_,cycle},l}$, and $P_{\mathrm{\_,cycle}}$ is challenging due to the overall complexity of computing systems. For instance, $N_{\mathrm{\_,cycle,}l}$ varies across microprocessors when processing the same amount of data, and it does not have a linear dependency with data size [33]. Therefore, the values for $B_{\mathrm{T}_{l}}$ and $N_{\mathrm{\_,cycle}}$ cannot be chosen arbitrarily. Next, $P_{\mathrm{\_,cycle}}$ also depends on the microprocessor. However, in Fig. 3, we illustrate the energy consumption of $E_{\mathrm{DC}}$ with a selected set of values to provide an intuition on the contribution from each term. More specifically, we analyze a scenario where $N_{\mathrm{S}}=256$, and we consider two sets of parameters to compute transmitting and receiving energy, i.e. $E_{\mathrm{T}}$ and $E_{\mathrm{R}}$, alongside two different settings for calculating the computational energy consumption ($E_{\mathrm{C}}$). The results show that while transmission energy ($E_{\mathrm{T}}$) remains the dominant contributor in most cases, the computational energy consumption ($E_{\mathrm{C}}$) is not negligible and, in some configurations, constitutes a significant portion of the total energy consumption of data collection. This is particularly the case when the hardware is less optimized for the tasks [33], leading to higher cycles per bit ($N_{\mathrm{\_,cycle},l}$) and power per cycle ($P_{\mathrm{\_,cycle}}$). For example, when $N_{\mathrm{\_,cycle}}=100$ and $P_{\mathrm{\_,cycle}}=2\times 10^{-10}$, $E_{\mathrm{C}}$ is $1.51\times 10^{-2}\penalty 10000\ [J]$, which represents approximately $10\%$ and $95\%$ of the total energy cost in two different PHY configurations. On the other hand, when $N_{\mathrm{\_,cycle},l}=50$ and $P_{\mathrm{\_,cycle}}=10^{-10}$, $E_{\mathrm{C}}$ is $6.61\times 10^{-3}\penalty 10000\ [J]$, contributing to approximately $5\%$ and $90\%$ of the total energy consumption in data collection component.
Figure 4: Proposed $E_{\mathrm{DC}}$ comparison with real-world measurement data.
Table III: Comparison of EE when transmitting $0.1$ Mbits
Tech EE w/ $E_{\mathrm{DC}}$ (Mbit/J) EE (Mbit/J)
Wi-Fi (8 UEs) 17.07 17.42
Wi-Fi (1 UE) 41.31 41.32
5G 1.63 1.65
Empirical Validation
To further validate our approach in a real-world setting, empirical measurements from Wi-Fi and 5G networks are considered. Specifically, in [34], the authors conducted rigorous measurements of various parameters, such as orthogonal frequency-division multiple access (OFDMA) throughput, latency, and power consumption achieved by Wi-Fi 6 technology, i.e., IEEE 802.11ax, using commercial devices. For 5G measurements, we consider the experiment conducted in [35], where a mobile user equipment (UE) is connected to a commercial 5G modem.
The results are shown in Fig. 4. The proposed $E_{\mathrm{DC}}$ consistently serves as a tight upper bound across both technologies, and hence corresponds to a lower bound on achievable EE (bits per Joule). For example, under the same bit window size of $0.75$ to $1.2\times 10^{5}$ [$b$] the measured energy consumption ($E_{\mathrm{COMM,\_}}$) of 5G (highlighted in blue), Wi-Fi 6 in a multi-access scenario ($8$ UEs, highlighted in red), and single UE (highlighted in orange) are $8.4$ to $12.4$ [mJmJ], $4.1$ to $7.2$ [mJmJ], and $1.4$ to $2.75$ [mJmJ], respectively. The proposed approach to the corresponding cases are $9.38$ to $15.8$ [mJmJ], $6.03$ to $8.63$ [mJmJ], $1.82$ to $3.41$ [mJmJ]. Moreover, the comparison of the EE when transmitting $0.1$ Mbits of data can be seen in Table III. As discussed, the proposed approach serves as a tight lower bound for EE. In other words, if we put the energy consumption of the computational part into consideration, this data-manipulating component will be less energy efficient than only considering physical layer measurement.
The methodology used for obtaining these results is as follows. We aggregate and process three datasets that are collected in real-world measurements, corresponding to uplink with $8$ UEs (1 mobile UE and 7 Personal Computers with Wi-Fi), uplink with a single mobile UE (Wi-Fi), and single UE in a commercial 5G network. For the Wi-Fi scenarios, we extract device-side transmission energy and device specifications from measurements to parameterize the UE terms in Eq. (3); for the access point (AP) (ASUS RT-AX58U), we use hardware specifications to obtain the receiving and computing terms in Eq. (3). To model the OSI‐stack overhead, we use published cryptography and networking guidance. On mobile UE (Galaxy S10, ARMv8 with cryptography extensions) [36], we assign $0.3$ [$c/b$] from transport to application layer (L4 to L7) and $0.1$ [$c/b$] at the medium access control (MAC) and Network layer (L2 and L3). These numbers are consistent with Advanced Encryption Standard (AES)-based encryption throughput on ARM-based architectures with hardware acceleration reported in OpenSSL microbenchmarks [37]. On PCs (Intel AX210 NICs) and AP, link-layer security and MAC functions are Network Interface Card (NIC)-offloaded. Therefore, we assume $0.02$ [$c/b$] (application), $0.01$ [$c/b$] (TCP/IP protocol), and $0.005$ [$c/b$] (MAC), aligned with Intel AES documentation and NIC offload literature [38, 39]. Guided by [40], we set $0.3$ [$nJ/c$] and $0.1$ [$nJ/c$] on ARM and Intel-based systems, respectively. For 5G measurement, we can obtain the receiving power and the receiving data size. It is mentioned in [41] that the pathloss follows the 3GPP UMa model, so we can then obtain the transmitting power. To account for protocol stack overhead, we set the overhead for L2 to L6 as $0.2$ and L7 as $0.1$. Since the setup of UE and AP is similar to the Wi-Fi scenario (same UE and Intel-based processor), the computing parameters are assumed to be the same as in the Wi-Fi scenario. These parameters help to anchor the derived $E_{\mathrm{DC}}$ to realistic hardware capabilities at practical operating points.
@@ V Energy Cost of Data Preprocessing
At the server side, data is first stored on a hard drive, subsequently preprocessed, and finally used for the ML model development or inference. This section focuses on preprocessing, while inference is addressed in Section VII.
Storage is generally energy-efficient, with relatively negligible energy costs. However, it can become significant when handling large volumes of incoming data or when utilizing virtualized hosted storage with encryption in an edge-cloud environment [42]. In the following case study, we consider the dataset with $N_{\mathrm{S}}$ samples collected via the wireless network, as depicted in Fig. 1 and discussed in Section IV. We also consider data preprocessing consisting of data cleaning, and data transformation steps, while leaving other more complex approaches, such as feature engineering, for further extensions of the study and simulation tool.
In the data cleaning step, the task of preprocessing is to identify and remove invalid samples from a list or array, which incurs zero floating-point operations (FLOPs) [^3] FLOPs are often used to enable theoretical estimation of computational complexity. Alternative approaches would involve direct energy measurements through interfaces such as PAPI or tools such as Scaphandre., as no floating-point additions or multiplications are performed, only comparisons and memory operations. During the data transformation step, preprocessing involves determining the complexity of data transformation ($M_{\mathrm{DS}}$). We provide three examples, i.e., min-max scaling, normalization, and Gramian Angular Difference Field (GADF) process, which represent typical approaches for data transformation.
@@
Standard score, also known as z-score normalization, represents a more computationally intensive preprocessing technique that involves three distinct computational steps. First, calculating the population mean requires $N_{\mathrm{S}}I_{\mathrm{S}}$ FLOPs, as we sum all values and divide by the total number of samples. Second, computing the population standard deviation demands $3N_{\mathrm{S}}I_{\mathrm{S}}+1$ FLOPs, which encompasses calculating squared differences from the mean, their sum, and the final square root operation. Finally, the standardization transformation necessitates $2N_{\mathrm{S}}I_{\mathrm{S}}$ FLOPs, as each data point must be centered by subtracting the mean and divided by the standard deviation for scaling. The total computational cost for normalization, expressed in FLOPs ($M^{\mathrm{norm}}_{\mathrm{DS}}$), is therefore:
@@
To showcase an even more computationally expensive preprocessing operation, we implemented a calculator to compute the energy cost of the GADF transformation [43]. This transformation converts time series data into an image representation that captures temporal correlations between each pair of values in the time series. The GADF computation consists of several sequential steps, each contributing to the overall computational complexity and, consequently, the energy consumption. The first step requires scaling the time series data using min-max scaling, which ensures that values fall within the [-1, 1] interval. Following the scaling, the data undergoes conversion to polar coordinates, where the radius is calculated using the time stamp values, and the angular coordinates are computed through an inverse cosine operation. The final step involves constructing the GADF matrix through pairwise angle differences, resulting in an $I_{\mathrm{S}}\times I_{\mathrm{S}}$ matrix where $I_{\mathrm{S}}$ represents the length of the input time series.
The computational complexity of GADF can be broken down into its constituent operations. The initial min-max scaling requires $2N_{\mathrm{S}}I_{\mathrm{S}}+1$ FLOPs as previously discussed. However, in this case we have to multiply the number of samples with the length of the input time series ($I_{\mathrm{S}}$) as we are dealing with a series of values in a single sample increasing the FLOPs. The polar coordinate transformation and the construction of the GADF matrix requires $(5I_{\mathrm{S}}+I_{\mathrm{S}}^{2})N_{\mathrm{S}}$ FLOPs . Therefore, the total number of FLOPs ($M^{\mathrm{GADF}}_{\mathrm{DS}}$) for the GADF transformation can be expressed as:
@@
This quadratic computational complexity makes GADF significantly more energy-intensive compared to simpler preprocessing operations like standardization or min-max scaling, particularly for longer time series. As a result, the total number of FLOPs for preprocessing can be expressed as:
@@
With the computing complexity in FLOPs formalized, we can obtain total energy consumption ($E_{\mathrm{pre}}$) of the preprocessing ($E_{\mathrm{pre}}$) by utilizing the power consumption of the processing unit ($P_{\mathrm{pre}}$) and the executing time ($T_{\mathrm{pre}}$):
@@
Numerical Validation
To illustrate the energy cost in various $N_{\mathrm{S}}$ and $I_{\mathrm{S}}$ configurations, consider the preprocessing energy cost $E_{\mathrm{pre}}$ on a CPU with $M_{\mathrm{PU}}=10\penalty 10000\ \text{GFLOPs/s}$ and power consumption of $P_{\mathrm{pre}}=140\penalty 10000\ \text{W}$, shown in Fig. 5 and Fig. 6. More specifically, Fig. 5 demonstrates how the energy consumption of preprocessing methods for a fixed number of samples ($N_{\mathrm{S}}=256$) increases with the sample size ($I_{\mathrm{S}}$). This trend is consistent across min-max scaling, normalization, and GADF, with GADF exhibiting the most pronounced energy growth due to its quadratic dependency on $I_{\mathrm{S}}$, whereas, the energy cost of the other two techniques grows linearly. In addition, Fig. 6 focuses on the energy consumption of the preprocessing techniques while maintaining a constant total amount of data ($N_{\mathrm{S}}I_{\mathrm{S}}=100,000$). This constraint ensures that as the sample size ($I_{\mathrm{S}}$) increases, the number of data samples ($N_{\mathrm{S}}$) decreases proportionally. Consequently, the energy consumption for min-max scaling and normalization remains constant across different configurations, as their computational complexity scales linearly with $N_{\mathrm{S}}I_{\mathrm{S}}$. In contrast, GADF exhibits a quadratic (with a linear component) growth in energy consumption due to the $I^{2}_{\mathrm{S}}$ term in its computational complexity, shown in Eq. (8), which is also confirmed by the $O\left(N_{\mathrm{S}}I^{2}_{\mathrm{S}}\right)$ shown in the gray dotted line. These results highlight the trade-offs in selecting a preprocessing method based on both the EE requirements and the specific input data characteristics required by the AI/ML technique.